# Estimating Detector Error Models on Google’s Willow

### v1

@@ IV Analysis of Google’s Willow Using DEMs
We sincerely appreciate Google releasing a large corpus of QEC data, complete with stim circuits, from experiments with their 72- and 105-qubit chips acharya2024quantumerrorcorrectionsurface. In this section, we apply the learning methods above to Google’s data and demonstrate their utility in exploratory data analysis. Some of the artifacts found by these methods are interesting in their own right, so we report them here.
@@ IV.2 Comparison of Estimated Excitation Rates
In addition the corresponding SI1000 DEM, each data-set in Google corpus for the $d=7$ surface code contains a DEM estimated by reinforcement learning (RL) on sub-patches of the $d=7$ surface code sivak2024optimizationdecoderpriorsaccurate. The SI1000 and RL-Prior DEMs provided by Google are structurally identical but differ in the excitation rates. Starting with this structure, we compare the rates estimated by Algorithm 3 to those learned by RL and those chosen a priori in SI1000. To enrich the comparison, we identify six hyperedge classes present in the temporal bulk of these DEMs:
@@
Figure 6: Cumulative distribution functions of hyperedge excitation rates. The six plots correspond to six classes of hyperedges discussed in the text. Within each plot, blue lines represent SI1000 parameters, orange lines represent parameters learned by Algorithm 3++ using syndromes from the 105-qubit chip, and green lines show parameters estimated by reinforcement learning as described in sivak2024optimizationdecoderpriorsaccurate.
@@
We quantify the effect of DEM priors on decoding performance by decoding a subset of Google’s data using two decoders: PyMatching Higgot_2025 and stim’s BP-OSD implementation gidney2021stim; roffe_decoding_2020, and three priors: Google’s RL prior, the SI1000 prior, and our estimated prior. At this time, we cannot use a DEM whose structure has been learned from syndromes as a prior for a decoder, because we lack an efficient means of associating a learned hyperedge with a change in the logical observable which the decoder requires. Bridging this gap is a subject for future work.
Google’s SI1000 prior is a simple, parameterized prior tailored to superconducting qubits, but not fitted specifically to a piece of hardware. It does not differ on a qubit-to-qubit basis. In contrast, Google’s RL prior was trained as in sivak2024optimizationdecoderpriorsaccurate and fine tuned on the provided 13-round “calibration data-set” taken directly from Google’s 105-qubit chip.
@@ IV.3 Divergences of DEMs from Syndromes
The Kullback-Leibler (KL) divergence kullback-leibler-divergence is a standard measure of disagreement between two probability distributions. Here, we adapt it to measuring divergence between syndromes and a DEM. In this scenario, the KL divergence of a DEM, $\mathcal{D}$, from a set of syndromes, $\mathbf{X}$, can be interpreted as the expected cost (in natural log-units) of encoding $\mathbf{X}$ using a code optimized for $\mathcal{D}$. The KL divergence is given by
@@ IV.5.1 Correlated Measurement Error
Although this analysis does not conclusively diagnose correlated measurement error, it does nominate a strong lead for further investigation and demonstrates the utility of a multi-stage exploratory analysis starting with pairwise correlations and exploiting various DEM estimation techniques for targeted interrogation of anomalies. The occurrence of these and other more exotic error classes may be reflected in future error budgets estimated in the manner of acharya2024quantumerrorcorrectionsurface.
@@ IV.5.2 High-Energy Events
To search for the signature of high-energy events acharya2024quantumerrorcorrectionsurface, we locate the shot in each sample with the largest shot-detector-event-fraction in raw and smoothed (over ten shots) where peaks in both traces are within five shots. This selection yields 30 candidates. [^3] We also undertook a broader selection allowing more than one candidate per sample, but found no more high-energy events. This secondary check was useful as part of our systematic uncertainty analyses. We proceed to zoom in, changing time from shots to rounds, then identify the maximum round-detector-event-fraction followed by an apparent exponential falloff, examined manually. This selection process is shown by example in Figure 12(a)-(c) with a spatial heat-map of qubit-detector-event-fraction in Figure 12(d) demonstrating this example event activity is not strictly correlated to the repetition code connectivity.
@@
Of the 30 candidates, we identified 23 as high-energy events based on their exponential decay. We overlay the selected 23 event windows of smoothed round-detector-event-fraction, aligned along their rising edges, in Figure 12(e). Note that we discovered approximately $4\times$ more high-energy events than acharya2024quantumerrorcorrectionsurface. We anticipate that our threshold is looser than that publication. Which threshold is appropriate for determining impact to QEC is a matter for future work.
@@ V Future Work
  • We suspect that techniques from log-linear analysis, e.g. those described in 10.1145/2939672.2932775, could improve the efficiency of structure learning.
@@ VI Conclusions
Arguably, the reinforcement learning employed in sivak2024optimizationdecoderpriorsaccurate belongs to a third class of DEM estimation algorithms, albeit one that serves a different purpose than the algorithms discussed here. On the one hand, because RL DEMs are trained with a decoder in the loop, they excel as priors for decoders used in logical memory experiments. On the other hand, when the goal is to understand device physics, the choice of decoder becomes an unwanted degree of freedom. As demonstrated with a $d=3$ surface code, RL DEMs do not exhibit the same fidelity to syndromes as DEMs estimated by decoder-free algorithms.
@@ Appendix A Proofs
Given $F(A,B)=G(A\cap B)H(A-B)$ with $F:\mathcal{P}([n])^{2}\to\mathbb{R}$ and $G,H:\mathcal{P}([n])\to\mathbb{R}$. Then for all $S,B\subseteq[n]$,
$\sum_{A\subseteq S}F(A,B)=\left[\sum_{C\subseteq S\cap B}G(C)\right]\times\left[\sum_{D\subseteq S-B}H(D)\right].$ (62)
@@
Recall - given any two sets $S,B$ - that $S=(S\cap B)\cup(S-B)$. Now for any $A\subseteq S$, there exists exactly one pair $(C,D)\in\mathcal{P}(S\cap B)\otimes\mathcal{P}(S-B)$ such that $A=C\cap D$. Specifically, $(C,D)=(A\cap B,A-B)$. Returning to the LHS of Eq. (62),
$\sum_{A\subseteq S}F(A,B)$ $=\sum_{A\subseteq S}G(A\cap B)H(A-B)$ (hypothesis)
$=\sum_{C\subseteq S\cap B}\sum_{D\subseteq S-B}G(C)H(D)$ (uniqueness)
$=\left[\sum_{C\subseteq S\cap B}G(C)\right]\times\left[\sum_{D\subseteq S-B}H(D)\right].$
@@
$=\sum_{B\subseteq[n]}(-1)^{|B|}\times\sum_{C\subseteq\{n+1\}}(-1)^{|C|}$ (Lemma 1)

### v2

@@ IV Analysis of Google’s Willow Using DEMs
We sincerely appreciate Google releasing a large corpus of QEC data, complete with stim circuits, from experiments with their 72- and 105-qubit chips AI and Collaborators (2024). In this section, we apply the learning methods above to Google’s data and demonstrate their utility in exploratory data analysis. Some of the artifacts found by these methods are interesting in their own right, so we report them here.
@@ IV.2 Comparison of Estimated Excitation Rates
In addition the corresponding SI1000 DEM, each data-set in Google corpus for the $d=7$ surface code contains a DEM estimated by reinforcement learning (RL) on sub-patches of the $d=7$ surface code Sivak et al. (2024). The SI1000 and RL-Prior DEMs provided by Google are structurally identical but differ in the excitation rates. Starting with this structure, we compare the rates estimated by Algorithm 3 to those learned by RL and those chosen a priori in SI1000. To enrich the comparison, we identify six hyperedge classes present in the temporal bulk of these DEMs:
@@
Figure 6: Cumulative distribution functions of hyperedge excitation rates. The six plots correspond to six classes of hyperedges discussed in the text. Within each plot, blue lines represent SI1000 parameters, orange lines represent parameters learned by Algorithm 3++ using syndromes from the 105-qubit chip, and green lines show parameters estimated by reinforcement learning as described in Sivak et al. (2024).
@@
We quantify the effect of DEM priors on decoding performance by decoding a subset of Google’s data using two decoders: PyMatching Higgott and Gidney (2025) and stim’s BP-OSD implementation Gidney (2021); Roffe et al. (2020), and three priors: Google’s RL prior, the SI1000 prior, and our estimated prior. At this time, we cannot use a DEM whose structure has been learned from syndromes as a prior for a decoder, because we lack an efficient means of associating a learned hyperedge with a change in the logical observable which the decoder requires. Bridging this gap is a subject for future work.
Google’s SI1000 prior is a simple, parameterized prior tailored to superconducting qubits, but not fitted specifically to a piece of hardware. It does not differ on a qubit-to-qubit basis. In contrast, Google’s RL prior was trained as in Sivak et al. (2024) and fine tuned on the provided 13-round “calibration data-set” taken directly from Google’s 105-qubit chip.
@@ IV.3 Divergences of DEMs from Syndromes
The Kullback-Leibler (KL) divergence Kullback and Leibler (1951) is a standard measure of disagreement between two probability distributions. Here, we adapt it to measuring divergence between syndromes and a DEM. In this scenario, the KL divergence of a DEM, $\mathcal{D}$, from a set of syndromes, $\mathbf{X}$, can be interpreted as the expected cost (in natural log-units) of encoding $\mathbf{X}$ using a code optimized for $\mathcal{D}$. The KL divergence is given by
@@ IV.5.1 Correlated Measurement Error
Although this analysis does not conclusively diagnose correlated measurement error, it does nominate a strong lead for further investigation and demonstrates the utility of a multi-stage exploratory analysis starting with pairwise correlations and exploiting various DEM estimation techniques for targeted interrogation of anomalies. The occurrence of these and other more exotic error classes may be reflected in future error budgets estimated in the manner of AI and Collaborators (2024).
@@ IV.5.2 High-Energy Events
To search for the signature of high-energy events AI and Collaborators (2024), we locate the shot in each sample with the largest shot-detector-event-fraction in raw and smoothed (over ten shots) where peaks in both traces are within five shots. This selection yields 30 candidates. [^3] We also undertook a broader selection allowing more than one candidate per sample, but found no more high-energy events. This secondary check was useful as part of our systematic uncertainty analyses. We proceed to zoom in, changing time from shots to rounds, then identify the maximum round-detector-event-fraction followed by an apparent exponential falloff, examined manually. This selection process is shown by example in Figure 12(a)-(c) with a spatial heat-map of qubit-detector-event-fraction in Figure 12(d) demonstrating this example event activity is not strictly correlated to the repetition code connectivity.
@@
Of the 30 candidates, we identified 23 as high-energy events based on their exponential decay. We overlay the selected 23 event windows of smoothed round-detector-event-fraction, aligned along their rising edges, in Figure 12(e). Note that we discovered approximately $4\times$ more high-energy events than AI and Collaborators (2024). We anticipate that our threshold is looser than that publication. Which threshold is appropriate for determining impact to QEC is a matter for future work.
@@ V Future Work
  • We suspect that techniques from log-linear analysis, e.g. those described in Webb and Petitjean (2016), could improve the efficiency of structure learning.
@@ VI Conclusions
Arguably, the reinforcement learning employed in Sivak et al. (2024) belongs to a third class of DEM estimation algorithms, albeit one that serves a different purpose than the algorithms discussed here. On the one hand, because RL DEMs are trained with a decoder in the loop, they excel as priors for decoders used in logical memory experiments. On the other hand, when the goal is to understand device physics, the choice of decoder becomes an unwanted degree of freedom. As demonstrated with a $d=3$ surface code, RL DEMs do not exhibit the same fidelity to syndromes as DEMs estimated by decoder-free algorithms.
@@ Appendix A Proofs
Let there be a set $S$ with subset $B$ and function $f:\mathcal{P}S\otimes\mathcal{P}S\to\mathbb{F}$ for arbitrary field $\mathbb{F}$. Then
$\sum_{A\subseteq S}f(A,B)=\left[\sum_{C\subseteq S\cap B}g(C)\right]\times\left[\sum_{D\subseteq S-B}h(D)\right].$
if and only if there exist functions $g,h:\mathcal{P}S\to\mathbb{F}$ such that $f(A,B)=g(A\cap B)h(A-B)$ for all $A\subseteq S$.
@@
Define the functions $d_{B}:\mathcal{P}S\to\mathcal{P}B\otimes\mathcal{P}(S-B)$ as $d_{B}(A)=(A\cap B,A-B)$ and $m_{B}:\mathcal{P}B\otimes\mathcal{P}(S-B)\to\mathcal{P}S$ as $m_{B}(C,D)=C\cup D$. Observe that $m_{B}(d_{B}(A))=A$ so both functions are bijective.
Observe then:
$\left[\sum_{C\subseteq S\cap B}g(C)\right]\left[\sum_{D\subseteq S-B}h(D)\right]=\sum_{C\subseteq S\cap B}\sum_{D\subseteq S-B}g(C)h(D)=\sum_{A\subseteq S}g(A\cap B)h(A-B)$
by field distributivity and the bijectivity of $d_{B}$ respectively. Now if the sum and product-of-sums are equal, we move forward to see that such $g,h$ exist. Meanwhile if we assume $f(A,B)=g(A\cap B)h(A-B)$ for all $A$, we move backward and find the sum and product-of-sums are equivalent. ∎
@@
$=\left[\sum_{B\subseteq[n]}(-1)^{|B|}\right]\left[\sum_{C\subseteq\{n+1\}}(-1)^{|C|}\right]$ (Lemma 1)